40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 11 04:08:53 2015, Jenne, Update, LSC, New Locking Paradigm? 
    Reply  Wed Feb 11 22:13:44 2015, Jenne, Update, LSC, New Locking Paradigm - LSC model changes 
       Reply  Thu Feb 12 01:43:09 2015, rana, Update, LSC, New Locking Paradigm - LSC model changes TF.gif
       Reply  Thu Feb 12 11:14:29 2015, Jenne, Update, LSC, New Locking Paradigm - Loop-gebra ALS_REFL_comboLockingCartoon_11Feb2015.PDFALS_REFL_comboLocking_11Feb2015.zip
          Reply  Thu Feb 12 11:59:58 2015, Koji, Update, LSC, New Locking Paradigm - Loop-gebra 
             Reply  Thu Feb 12 19:18:49 2015, Jenne, Update, LSC, New Locking Paradigm - Loop-gebra 
                Reply  Fri Feb 13 03:28:34 2015, rana, Update, LSC, New Locking Paradigm - Loop-gebra 
                   Reply  Sat Feb 14 20:20:24 2015, Jenne, Update, LSC, ALS fool cartoon ALSfool_LoopDiagram.pngALSfool_LoopDiagram.graffle.zip
                      Reply  Tue Feb 17 04:04:32 2015, Jenne, Update, LSC, ALS fool math OpenLoopGainComparison_16Feb2015.pngCancellationTFs_DecouplingIsArefl.png
                         Reply  Tue Feb 17 16:36:08 2015, Jenne, Update, LSC, ALS fool math ALSfool_LoopDiagram_17Feb2015.pngALS_REFL_comboLocking_16Feb2015.nb.zip
       Reply  Thu Feb 12 22:28:16 2015, Jenne, Update, LSC, New Locking Paradigm - LSC model changes, screens modified 
          Reply  Sat Mar 7 19:15:17 2015, Jenne, Update, LSC, Modified zero crossing triggering LatchLogic_6March2015.png
    Reply  Thu Feb 12 03:43:54 2015, ericq, Update, LSC, 3F PRMI at zero ALS CARM Feb12_negativeTR.png
Message ID: 11010     Entry time: Thu Feb 12 03:43:54 2015     In reply to: 11001
Author: ericq 
Type: Update 
Category: LSC 
Subject: 3F PRMI at zero ALS CARM 

I have been able to recover the ability to sit at zero CARM offset while the PRMI is locked on RELF33 and CARM/DARM are on ALS, effectively indefinitely. However, I feel like the transmon QPDs are not behaving ideally, because the reported arm powers freqently go negative as the interferometer is "buzzing" through resonance, so I'm not sure how useful they'll be as normalizing signals for REFL11. I tried tweaking the DARM offset to help the buildup, since ALS is only roughly centered on zero for both CARM and DARM, but didn't have much luck.

Example:

Turning off the whitening on the QPD segments seems to make everything saturate, so some thinking with daytime brain is in order.


How I got there:

It turns out triggering is more important than the phase margin story I had been telling myself. Also, I lost a lot of time to needing demod angle change in REFL33. Maybe I somehow caused this when I was all up on the LSC rack today?

We have previously put TRX and TRY triggering elements into the PRCL and MICH rows, to guard against temporary POP22 dips, because if arm powers are greater than 1, power recylcing is happening, so we should keep the loops engaged. However, since TRX and TRY are going negative when we buzz back and forth through the resonsnace, the trigger row sums to a negative value, and the PRMI loops give up. 

Instead, we can used the fortuitously unwhitened POPDC, which can serve the same function, and does not have the tendancy to go negative. Once I enabled this, I was able to just sit there as the IFO angrily buzzed at me. 

Here are my PRMI settings

REFL33 - Rotation 140.2 Degrees, -89.794 measured diff

PRCL = 1 x REFL33 I; G = -0.03; Acquire FMs 4,5; Trigger FMs 2, 9; Limit: 15k ; Acutate 1 x PRM

MICH = 1 x REFL33 Q, G= 3.0, Acquire FMs 4,5,8; Trigger FM 2, 3; Limit: 30k; Actuate -0.2625 x PRM + 0.5 x BS

Triggers = 1 x POP22 I + 0.1 * POPDC, 50 up 5 down


Just for kicks, here's a video of the buzzing as experienced in the control room

Attachment 1: Feb12_negativeTR.png  494 kB  | Show | Show all
ELOG V3.1.3-